### v1

# No Thing, Nothing: Highlighting Safety-Critical Classes for Robust LiDAR Semantic Segmentation in Adverse Weather

Abstract
Existing domain generalization methods for LiDAR semantic segmentation under adverse weather struggle to accurately predict “things” categories compared to “stuff” categories. In typical driving scenes, “things” categories can be dynamic and associated with higher collision risks, making them crucial for safe navigation and planning. Recognizing the importance of “things” categories, we identify their performance drop as a serious bottleneck in existing approaches. We observed that adverse weather induces degradation of semantic-level features and both corruption of local features, leading to a misprediction of “things” as “stuff”. To mitigate these corruptions, we suggests our method, NTN - segmeNt Things for No-accident. To address semantic-level feature corruption, we bind each point feature to its superclass, preventing the misprediction of things classes into visually dissimilar categories. Additionally, to enhance robustness against local corruption caused by adverse weather, we define each LiDAR beam as a local region and propose a regularization term that aligns the clean data with its corrupted counterpart in feature space. NTN achieves state-of-the-art performance with a +2.6 mIoU gain on the SemanticKITTI-to-SemanticSTF benchmark and +7.9 mIoU on the SemanticPOSS-to-SemanticSTF benchmark. Notably, NTN achieves a +4.8 and +7.9 mIoU improvement on “things” classes, respectively, highlighting its effectiveness.
1 Introduction
Figure 1: Previous methods struggle to accurately predict the things category in adverse weather. Our approach overcomes this limitation with improved performance for things category.
Figure 2: (a) Comparison of things classes in SemanticKITTI [3] (clean) and SemanticSTF [28] (corrupted) datasets. (b) Examples of things classes misclassified as stuff due to weather corruption. The prediction results were obtained with SJ+LPD [20].
LiDAR Semantic Segmentation (LSS) involves predicting semantic labels directly for each 3D point, an important building block for autonomous driving. However, LSS models exhibit a significant performance drop under adverse weather conditions such as fog, rain, and snow [28, 14, 20, 31, 10]. Since safety is the highest priority in autonomous driving, it is essential to develop robust models that can maintain high accuracy regardless of weather conditions.
To mitigate this issue, existing research on improving robustness under adverse weather conditions can be broadly categorized into domain generalization and domain adaptation methods. Domain adaptation methods [31] utilize real weather training data to train LSS models to be robust against weather corruption. In contrast, domain generalization methods aim to create models robust to weather corruption only using training data under clean weather. Typically, these studies either improve training mechanisms [28, 10] or generate LiDAR data through weather simulations [20, 8, 9]. Domain generalization in adverse weather is particularly important, as real weather varies in type and severity and occurs infrequently, making data collection challenging. In this paper, we focus on the domain generalization problem.
However, existing Domain Generalization methods for LiDAR Segmentation under Adverse Weather (DGLS-AW) show significant performance variations across semantic classes. In driving scenarios, semantic categories can be divided into stuff and things. Stuff shows continuous distributions and is static such as road and terrain. Meanwhile, things consist of distinct and dynamic instances like pedestrians, bicycles, cars, and trucks. As shown in Fig. 1, the segmentation models generalized for adverse weather [28, 14, 20, 31, 10] exhibit significantly lower prediction performance for things compared to stuff. Even the state-of-the-art model SJ+LPD [20] shows a significant performance gap, with things mIoU averaging 11.8 lower than stuff. This performance gap poses serious challenges for autonomous driving. Failing to accurately detect objects like pedestrians or vehicles can lead to critical incidents, as it hinders the ability to react quickly to their movements. On the other hand, missing stuff category generally has a lower impact on immediate accident risk.  [^1] It is well-known that misclassifying a things category as a stuff category can be fatal, as it fails to account for the potential movement of these objects. It was revealed that an accident that occurred in May 2016 while using a driving assistance system was caused by incorrectly predicting the white part of the trailer as the sky [11, 1]. For these reasons, accurately detecting things is a critical mission considering driving applications.
Our study focuses on mitigating performance limitations for the things classes in DGLS-AW. For that, we first identify key factors contributing to the performance gap and propose a solution to address each challenge: weather-induced corruption confuses the LSS model at both (1) semantic-level and (2) local-level features separately, especially for things classes.
First, adverse weather causes severe deformation, particularly affecting things classes, making their shapes unrecognizable and hindering the learning of semantic-level features. As shown in Fig. 2 (a), stuff classes (building, vegetation, and road) are large enough to cover the entire scene, therefore their geometric features remain stable regardless of whether the data is clean or corrupted. In contrast, things classes show surprisingly different appearances in clean and corrupted data.
This difference arises because things classes are generally small and dispersed, leaving them vulnerable to missing points that can distort their structure. For example, person class, belonging to the things category, have dense and clear shapes under clean weather, as shown on the far left of the figure. Consequently, under adverse weather conditions, the person and car points are fewer in number and have obscure shapes. We analyze that weather-induced corruption blurs the boundary between things and stuff, leading to frequent misclassification. Even the state-of-the-art method, SJ+LPD [20], often confused between things and stuff as shown in Fig. 2 (b).
To tackle this, we propose a Feature Binder (FB) that groups semantically similar classes into superclasses. By binding each point feature to its superclass, FB ensures that even in the presence of corruption, things classes can learn and retain their semantic information. This approach effectively reduces misclassification between things and stuff.
Second, the corruption leads to the loss of local-level features, which is especially critical for things classes due to their smaller size. Previous research states that adverse weather effects like point drops affect each laser beam independently, causing local feature corruption [8, 9, 20]. Since things classes rely heavily on distinctive local features (e.g., point patterns, intensity) for accurate prediction, this corruption is particularly detrimental to them. Missing points remove essential training signals from these regions, preventing effective feature learning for precise predictions.
To mitigate this issue, we introduce Beam-wise Feature Distillation (BFD), which aligns the features of corrupted data with those of clean data. By matching corrupted points with their clean counterparts, we provide additional information to the things classes. This addition allows the model to robustly predict things even under severe corruption, as the dense points from the clean data compensate for the missing information in the corrupted data.
We propose NTN (segmeNt Things for No-accident), a novel approach that integrates FB and BFD to tackle semantic- and local-level feature corruption in DGLSS-AW. NTN significantly improves the robustness of LSS models against adverse weather conditions, with particular benefits for safety-critical things classes. This approach achieves state-of-the-art performance on both the SemanticKITTI-to-SemanticSTF and SemanticPOSS-to-SemanticSTF benchmarks, with overall mIoU improvements of +2.6 and +7.9 for each, respectively. Notably, NTN demonstrates a +4.8 and +7.9 boost in things classes on these benchmarks, significantly increasing accuracy for highly mobile or structurally complex objects such as person, bicycle, and motorcyclist. These results show NTN’s ability to maintain precise segmentation of risk-sensitive objects in adverse weather scenarios.
In summary, our contributions are as follows:
  • 1.
    To our knowledge, we are the first to identify a significant performance degradation in safety-critical things classes under adverse weather conditions. We found that damage to both semantic- and local-level features caused by weather corruption is a key factor in performance degradation, particularly for things classes.
  • 2.
    To address these issues, we propose segmeNt Things for No-accident, which integrates two novel modules; Feature Binder and Beam-wise Feature Distillation.
  • 3.
    Our approach achieves state-of-the-art results on the SemanticKITTI-to-SemanticSTF and SemanticPOSS-to-SemanticSTF benchmarks, with significant gains of+4.8 and +7.9 in things classes performance, respectively.
2 Related Works
2.1 LiDAR Semantic Segmentation
LSS methods can be broadly categorized into (1) point-based, (2) projection-based, and (3) voxel-based approaches based on their input representations.
Point-based methods [21, 24, 30] process 3D points directly as input. PointNet [21] employs sequential MLPs for directly merging local features of points into global feature vector. KPConv [24] defines 3D kernel points and performs direct convolution with input points. Building upon this approach, KPConvX [25] introduces attention mechanisms within the kernel points, thereby enhancing the model capacity without significantly increasing the size of the 3D convolutional network. The Point-Mixer [4] adapts the MLP-Mixer [26] for point cloud processing. The Point Transformer [30] leverages a transformer architecture to compute query points within each local neighborhood found by k-nearest neighbors. While effective, these methods incur high computational costs due to processing large-scale raw LiDAR data. This is why point-based methods are not suitable for outdoor LSS.
Projection-based methods [17, 2, 13] project LiDAR points onto a 2D plane, allowing semantic segmentation to leverage architectures originally developed for 2D image tasks. RangeNet++ [17] was the first to address LSS using range images via spherical projection, adapting 2D convolutional neural networks for this purpose. RangeViT [2] applies a pre-trained ViT model on 2D images, demonstrating effective use of 2D image pre-training for range images. RangeFormer [13] introduces ”RangeAug,” which enhances 2D-projected range images by generating multiple inputs to boost model performance. Projection-based approaches offer fast inference but are limited by information loss during projection.
Voxel-based methods [5, 32, 15] divide 3D space into voxel grids for efficient computation. MinkowskiNet [5] voxelizes LiDAR data into cubic grids, using sparse convolutions. SPVCNN [23] builds on MinkowskiNet with an added point-wise MLP. Cylinder3D [32] uses cylindrical partitions, addressing LiDAR density variations with distance. SphereFormer [15] applies radial windows in voxelization and a transformer to aggregate long-range information. Voxel-based methods offer a balance of inference speed and segmentation accuracy.
2.2 LiDAR under Adverse Weather Conditions
Recently, LiDAR perception research in adverse weather has gained attention for its practical significance in autonomous driving, focusing on (1) weather simulation and (2) task-agnostic domain generalization methods.
Hahner et al. [9, 8] use physical modeling to create synthetic datasets for specific weather conditions, integrating them into training. UniMix [31] extends these works for adapting LiDAR segmentation models for adverse weather using these simulations. SJ+LPD [20] deviate from physical modeling, incorporating weather-induced corruption via point dropping and jittering to enhance robustness.
Recent task-agnostic approaches [14, 28, 10] employ general machine learning techniques, such as teacher-student frameworks and prototyping, to enhance robustness. Kong et al. [14] introduced random dropout in the teacher-student architecture to build noise-resistant models. PointDR [28] generates a prototype in the clean weather branch and multiplies corrupted branch features by this prototype matrix, utilizing clean data insights. DGUIL [10] addresses point uncertainty under domain shifts by by improving point features across varying distributions in uncertain point clouds.
However, existing weather simulation and domain generalization methods focus on replicating weather conditions and do not address precise recognition of key things classes under weather-induced corruption in driving environments. Building on SJ+LPD [20], our work proposes methods to enhance recognition of things classes.
Figure 3: Confusion matrices for (a) superclasses and (b) things and stuff. It shows a significant performance gap between things and stuff; things are often misclassified as stuff, whereas the reverse misclassification is rare.
3 Methods
Figure 4: (a) Framework of NTN. NTN builds upon SJ+LPD [20] and enhances the performance of the LSS model on the things category through Feature Binding (FB) and Beam-based Feature Distillation (BFD). (b) FB stores prototypes obtained from clean data in a memory bank and encourages features derived from augmented data to align with these prototypes. (c) BFD imposes constraints to ensure that features from clean data and those obtained from augmented data match within a manually defined local region, referred to as a beam.
3.1 Motivation
In this section, we aim to understand why existing LSS methods in adverse weather underperform on the things classes. To accomplish this, we analyze the impact of weather-induced corruption in adverse weather on 3D points. The findings in Sec. 1 can be summarized as follows: (1) Weather corruption degrades the semantics of things classes. (2) Local-level feature corruption causes significant missing points and leads to information loss.
As shown in Fig. 3, even the current state-of-the-art DGLS-AW method, SJ+LPD [20], misclassifies 22.66% of person predictions as natural, illustrating the tendency for things classes’ predictions to shift toward stuff. This misclassification occurs because the model struggles to maintain clear boundaries between visually distinct classes when the input features are corrupted. To address this, we constrain features from the LSS model to remain close within visually similar classes. This approach preserves semantic alignment and mitigates shifts between things and stuff, even under corruption-induced semantic degradation.
Also, as shown in Fig. 2 (a), significant missing points cause a loss of learning signals during training. This loss is particularly harmful to things classes. It is because things classes, with fewer points and smaller sizes, heavily rely on these learning signals. Specifically, in SemanticKITTI [3], things classes account for only 12.3% of the total points. The limited points in the things classes become even scarcer when missing, resulting the poor performance in the things category.
By focusing on these key points, we address the degraded performance of the things category in existing DGLS-AW methods. We adopt the existing adverse weather LSS, SJ+LPD [20] as a baseline, which achieves state-of-the-art performance through specialized weather simulation augmentations. Importantly, our proposed framework is architecture-independent and applicable to various baseline models beyond SJ+LPD, as it introduces new training schemes.
3.2 Weather-corrupted LiDAR Data Generation
Existing LiDAR weather simulation methods [8, 9, 20] simulate corrupted data $X^{aug}$ and use them as additional training data to achieve generalized performance under adverse weather. Hahner et al. [9, 8] employ physical equations to synthesize weather-specific LiDAR data for training. SJ+LPD [20], which we set as a baseline method, generates output points $\hat{X}$ by applying selective jittering to the original clean data. Subsequently, $\hat{X}$ is input into the Learnable Point Drop module $Q(\cdot)$ to produce the final augmented data $X^{aug}$. In subsequent methods, we utilize $\hat{X}$ as clean data (i.e., the original LiDAR points) and $X^{aug}$ as augmented data (i.e., the simulated or augmented counterpart).
Category Superclass SemanticKITTI SemanticPOSS Things Vehicle car, bicycle, motorcycle, truck, bus car, bicycle Person person, bicyclist, motorcyclist pedestrian Traffic Element fence, pole, traffic-sign - Stuff Pavement road, parking, sidewalk drivable-surface Natural other-ground, vegetation, terrain walkable Structure building, trunk -
Table 1: Superclasses of each category and its constituent classes for SemanticKITTI and SemanticPOSS datasets.
3.3 Feature Binding Loss
Building on the analysis in Sec. 3.1, we first introduce a concept of Feature Binding loss (FB). The purpose of FB loss is to “bind” the output features of LSS model, preventing misprediction of things objects as stuff when local corruption severely affects semantic-level features.
To achieve this, we create prototypes for superclasses by grouping visually similar pre-defined classes during LSS training. Tab. S10 summarizes the superclass categories and their corresponding classes. Six superclasses are defined for SemanticKITTI [3], while four superclasses are defined for SemanticPOSS [19]. Following the method in  [12], we associated each class with its respective superclass. Classes with significantly different shapes, such as trunk and vegetation are assigned to separate superclasses. Additionally, classes previously treated as background, like a motorcyclist, were reassigned to the person class.
We then constrain the output features of the LSS model for augmented data $x^{aug}$ to remain close to these superclass prototypes. This approach prevents severe misprediction into visually dissimilar stuff classes, even if the semantic features of input points are significantly degraded by weather corruption.
Specifically, we compute the prototype for each class $c_{i}$ as the mean feature of its constituent classes and store it in the memory bank $\mathcal{M}$.
$P_{\mathcal{M}}^{c_{i}}=\frac{1}{|\hat{X}_{c_{i}}|}\sum_{\hat{x}_{c_{i}}\in\hat{X}_{c_{i}}}G(F(\hat{x}_{c_{i}})).$ (1)
$F$ is the LSS model, $G$ is a projector head with MLP, and $|\hat{X}_{c_{i}}|$ denotes the number of points labeled as $c_{i}$ in the entire class set $\hat{X}_{\mathcal{C}}$. These class-wise prototypes are updated using an exponential moving average during each training iteration. Subsequently, the superclass prototype $P^{\mathcal{S}_{n}}_{\mathcal{M}}$ is defined using the class-wise prototypes belonging to each superclass $\mathcal{S}_{n}$ as follows:
$P^{\mathcal{S}_{n}}_{\mathcal{M}}=\frac{1}{|\mathcal{S}_{n}|}\sum_{c_{i}\in\mathcal{S}_{n}}P^{c_{i}}_{\mathcal{M}}.$ (2)
where $|\mathcal{S}_{n}|$ represents the number of classes in superclass $\mathcal{S}_{n}$. We then fit the feature output from augmented data to each superclass prototype to compute the FB loss as follows:
$\mathcal{L}_{FB}=\left\|\frac{1}{|X_{c}^{aug}|}\sum_{x_{c}^{aug}\in X_{c}^{aug}}G(F(x_{c}^{aug}))-\frac{1}{|\mathcal{S}|}\sum_{\mathcal{S}_{n}\in\mathcal{S}}P_{\mathcal{M}}^{\mathcal{S}_{n}}\right\|^{2}.$ (3)
This approach enables the LSS model to bind semantic features to their appropriate superclass by creating prototypes of visually similar, pre-defined classes. This method enables the model to prevent misprediction of things as stuff, even when handling corrupted inputs.
Figure 5: Illustration of (a) Point-wise Feature Distillation and (b) Beam-wise Feature Distillation.
3.4 Beam-wise Feature Distillation
Method car bi.cle mt.cle truck oth-v. pers. bi.clst mt.clst road parki. sidew. othe.g. build. fence veget. trunk terra. pole traf. mIoU Oracle 89.4 42.1 0.0 59.9 61.2 69.6 39.0 0.0 82.2 21.5 58.2 45.6 86.1 63.6 80.2 52.0 77.6 50.1 61.7 54.7 Source-only 55.9 0.0 0.2 1.9 10.9 10.3 6.0 0.0 61.2 10.9 32.0 0.0 67.9 41.6 49.8 27.9 40.8 29.6 17.5 24.4 Dropout [22] 62.1 0.0 15.5 3.0 11.5 5.4 2.0 0.0 58.4 12.8 26.7 1.1 72.1 43.6 52.9 34.2 43.5 28.4 15.5 25.7 Perturbation 74.4 0.0 0.0 23.3 0.6 19.7 0.0 0.0 60.3 10.8 33.9 0.7 72.0 45.2 58.7 17.5 42.4 22.1 9.7 25.9 PolarMix [27] 57.8 1.8 3.6 3.7 26.5 3.7 26.5 0.0 65.7 2.9 35.9 48.7 71.0 48.7 53.8 20.5 45.4 25.9 15.8 26.5 MMD [16] 63.6 0.0 2.6 17.4 11.4 28.1 0.0 0.0 67.0 14.1 37.6 41.2 67.1 41.2 57.1 27.4 47.9 28.2 16.2 26.9 PCL [29] 65.9 0.0 0.4 17.3 8.4 8.4 8.4 8.4 59.6 12.0 35.3 63.1 74.0 47.5 60.7 15.8 48.9 26.1 27.5 26.4 PointDR [28] 67.3 0.0 4.5 19.9 18.8 2.7 0.0 0.0 62.6 12.9 38.6 43.8 73.3 43.8 56.4 32.2 45.7 28.7 27.4 28.6 DGLSS [12] 72.6 0.1 11.7 29.4 13.7 48.3 0.5 21.2 65.0 20.2 38.3 3.8 78.9 51.8 57.0 36.4 47.0 26.9 34.9 34.6 UniMix [31] 82.7 6.6 8.6 4.5 19.9 35.5 15.1 15.5 55.8 10.2 36.5 40.1 72.8 40.1 49.1 33.4 34.9 23.5 33.5 31.4 DGUIL [10] 77.9 10.6 19.1 26.0 9.7 46.3 6.0 9.3 69.1 18.0 38.6 9.4 73.3 51.2 60.8 30.9 50.8 31.8 22.3 35.5 SJ+LPD [20] 83.1 1.2 17.2 30.5 18.4 47.5 10.7 18.8 64.0 15.9 38.7 4.6 77.4 50.8 59.7 37.2 47.7 31.1 35.8 36.3 NTN (Ours) 83.3 3.7 31.3 36.2 18.2 53.3 6.8 55.9 67.2 18.1 37.2 5.4 72.1 41.8 58.0 36.0 46.0 28.2 39.8 38.9 $\uparrow$ to SJ+LPD (+0.2) (+2.5) (+14.1) (+5.7) (-0.2) (+5.8) (-3.9) (+37.1) (+3.2) (+2.2) (-1.5) (+0.8) (-5.3) (-9.0) (-1.7) (-1.2) (-1.7) (+2.9) (+4.0) (+2.6)
Table 2: Comparison of methods on the SemanticKITTI-to-SemanticSTF benchmark. Performance improvements of our method over the current state-of-the-art model, SJ+LPD, are shown in the last row, with red text indicating things classes increments and green text for stuff.
This section presents a method to generate additional learning signals for things classes, which are particularly vulnerable to weather corruption due to their limited representation in the dataset. Weather corruption primarily affects localized regions, causing significant point loss and disrupting learning signals. Without sufficient learning signals, the model struggles to capture the necessary features for these classes, leading to degraded performance. To address this issue, we design an additional loss that also operates within localized areas, compensating for the loss of learning signals.
To define the local regions, we slice the LiDAR scan into each LiDAR beam. A beam corresponds to the output response of the laser beam that the LiDAR sensor fires at regular pitch angle intervals. Each beam scans the things object into upper and lower local regions. Using these beams, we define a loss ensuring that the average of the clean and augmented data is identical for each beam, exclusively for the things classes.
Specifically, we first apply off-the-shelf beam separation algorithms [8, 9, 14] for LiDAR data to identify the LiDAR beam source of the points corresponding to each voxel. Then, for each things classes and each LiDAR beam, we apply an MSE loss between the feature of clean data and that of augmented data.
$\mathcal{F}_{mean}^{c,b}=\frac{1}{N_{c,b}}\sum_{i=1}^{N_{c,b}}F(\hat{x}_{i}),\quad\mathcal{F}_{mean}^{{}^{\prime}c,b}=\frac{1}{N_{c,b}^{{}^{\prime}}}\sum_{i=1}^{N_{c,b}^{{}^{\prime}}}F(x_{i}^{aug}).$ (4)
$\mathcal{L}_{BFD}=\sum_{c\in\mathcal{S}_{\text{things}}}\sum_{b\in\mathcal{B}}\left\|\mathcal{F}_{mean}^{c,b}-\mathcal{F}_{mean}^{{}^{\prime}c,b}\right\|^{2}.$ (5)
$\mathcal{S}_{\text{things}}$ denote the things classes, and $c$ a specific class. Let $\mathcal{B}$ represent all LiDAR beams, and $b$ a specific beam. $N_{c,b}$ is the number of points for class $c$ and beam $b$. $\mathcal{L}_{BFD}$ minimizes the difference between the average features of clean data, $\mathcal{F}_{mean}^{c,b}$, and that of augmented data, $\mathcal{F}_{mean}^{{}^{\prime}c,b}$, for each beam, specifically for the things classes, as shown in Fig. 5 (b).
Our proposed beam-wise feature matching focuses on local regions, enabling learning from all points in $\hat{X}$ and leveraging contextual information around each beam unlike point-wise feature distillation in Fig. 5 (a). Points-wise feature distillation matches individual point features for distillation without averaging. In this case, no distillation loss can be computed for points dropped from $\hat{X}$ when obtaining $X^{aug}$. Furthermore, surrounding points contribute no learning information, limiting the clean data branch to learning only from a subset of points in the augmented data branch. In contrast, our beam-wise feature matching overcomes these limitations by effectively handling local feature corruptions in things category objects, both before and after weather-induced corruption. Therefore, BFD is a effective solution for handling local feature corruption in things category objects, both before and after weather-induced corruption.
3.5 Overall Pipeline
The overall training pipeline is illustrated in Fig. 4. The process for obtaining clean and augmented data follows SJ+LPD [20]. In SJ+LPD, Selective Jittering (SJ) is applied to obtain clean data $\hat{x}_{i}$. The loss and entropy from this data are used as the state for the Learnable Point Drop (LPD). LPD then drops points, and the LSS model outputs features for the augmented data $x_{i}^{aug}$. Thus, we obtain the output feature of clean data $F(\hat{x}_{i})$ and that of augmented data $F(x_{i}^{aug})$. These features are used to compute the FB loss and BFD loss, as described earlier. Additionally, we include the cross-entropy loss from both clean and augmented data. The overall loss is:
$\begin{split}\mathcal{L}_{total}=&\ \mathcal{L}_{clean}+\lambda_{aug}\mathcal{L}_{aug}\\ &+\lambda_{FB}\mathcal{L}_{FB}+\lambda_{BFD}\mathcal{L}_{BFD}.\end{split}$ (6)
where $\lambda_{\cdot}$ means the weight for each loss.
4 Experiments
4.1 Experimental Setup
Dataset. We utilize three datasets for autonomous driving to train and evaluate the generalization performance of the LSS model under adverse weather conditions. SemanticKITTI dataset uses a 64-beam LiDAR and provides 19 class labels. Following the official sequence split, we use scenes 00-07 and 09-10 for training, comprising 19,130 frames. The SemanticPOSS dataset uses a 40-beam LiDAR and focuses on campus scenes with a high density of cars, riders, and pedestrians. This dataset provides 14 class labels, and we use scenes 00-02 and 04-05 for training, comprising 2,488 frames. SemanticSTF dataset is collected under real-world adverse weather conditions, including rain, fog, and snow. This dataset follows the same 19 class labels used in the SemanticKITTI benchmark. We validate our model on the SemanticSTF validation set to evaluate robustness in adverse weather.
[…]
Evaluation Metrics. We adopt Intersection over Union (IoU) as the evaluation metric for each segmentation class and calculate the mean IoU (mIoU) across all classes.
Implementation Details. We adopted MinkowskiNet-18 architectures [5] as our baseline models. The training was conducted using a stochastic gradient descent (SGD) optimizer with a learning rate of 0.24 and a weight decay of 0.0001. The configurations for SJ+LPD are the same as the previous work [20]. All experiments were executed on four NVIDIA A6000 GPUs for 30 epochs with a batch size of 2. The total runtime for each experiment ranged from 6 to 8 hours. The loss function weights were set to $\lambda_{\text{aug}}=0.5$, $\lambda_{\text{SPA}}=0.5$, and $\lambda_{\text{BFD}}=0.1$.
Method Car Bicycle Person Road Walkable mIoU Source-only [28] 46.6 3.1 54.2 47.0 37.2 37.6 PointDR [28] 35.0 3.1 59.7 60.7 14.9 34.7 DGLSS [12] 60.0 3.1 38.6 54.0 40.1 39.2 SJ+LPD [20] 67.0 0.9 45.9 43.3 34.7 38.3 NTN (Ours) 65.5 5.8 61.4 54.6 43.6 46.2 $\uparrow$ to SJ+LPD (-1.5) (+4.9) (+15.5) (+11.3) (+8.9) (+7.9)
Table 3: Comparison of methods on the SemanticPOSS-to-SemanticSTF benchmark.
4.2 Main Results
SemanticKITTI to SemanticSTF. As shown in Table 2, our method achieves a state-of-the-art performance with a +2.6 mIoU improvement over the prior leading model, SJ+LPD [20]. Additionally, our approach shows a significant boost of +14.5 mIoU over the source-only model, the original MinkUnet. Furthermore, our method achieves notable improvements across most things classes, yielding gains of +2.5, +14.1, +5.7, +5.8, +37.1, +2.9, and +4.0 mIoU in bicycle, motorcycle, truck, person, motorcyclist, pole, and traffic sign classes, respectively. These results provide strong evidence that our method significantly enhances performance specifically for things classes, as intended. Notably, previous methods struggled to predict tail classes with small-sized, such as motorcycle and motorcyclist, accurately. It indicates that our method offers precise and reliable signals for identifying things classes under adverse conditions.
Prototype Feature Distillation Performance Matching Classes Things Stuff mIoU 31.4 43.2 36.3 Class-wise 31.6 42.2 36.1 Superclass 34.9 ($\uparrow$3.5) 42.5 38.1 ($\uparrow$1.8) Coarse 35.6 41.0 37.9 beam-wise all 34.0 43.3 37.9 point-wise things 33.1 41.7 36.7 beam-wise things 34.4 ($\uparrow$3.0) 44.0 38.4 ($\uparrow$2.1) Superclass beam-wise things 36.2 ($\uparrow$4.8) 42.5 38.9 ($\uparrow$2.6)
Table 4: Ablation study for different prototype and feature distillation settings. Coarse refers to the use of two prototypes divided into things and stuff for feature binding
SemanticPOSS to SemanticSTF. As shown in Table 3, our method establishes a new state-of-the-art on the SemanticPOSS-to-SemanticSTF benchmark, surpassing the performance of the previous model, SJ+LPD [20], by an impressive +7.9 mIoU. Importantly, in the person category, our method achieves a +15.5 mIoU gain over SJ+LPD. These results underscore the effectiveness of our approach in providing highly accurate segmentation signals for things classes under challenging adverse conditions. The SemanticPOSS dataset contains a higher proportion of things classes at 28.9%, compared to SemanticSTF, and is known for its diversity. Consequently, the model has already learned relatively rich features, resulting in less performance improvement for things classes compared to SemanticKITTI.
Figure 6: Qualitative results of our method on validation set of SemanticSTF. All models are trained on the train set of SemanticKITTI. Gray points indicate correct predictions, red points highlight errors, and ground-truth is shown with color-coded labels. Dashed circles highlight the predictions of things classes. Best viewed in color.
4.3 Ablataion Study
All ablation studies are conducted on the SemanticKITTI-to-SemanticSTF benchmark.
Feature Binding Loss. We conduct an ablation study to evaluate the impact of various prototype strategies of Feature Binding. We compare three approaches: assigning prototypes to each class individually, grouping semantically similar classes into superclasses, and broadly categorizing classes into coarse prototypes as stuff or things. As shown in Table 4, using superclass prototypes improves things classes performance by +3.3 mIoU compared to the class-wise method. This improvement indicates that our FB method effectively bounds things classes, which are prone to be misclassified as stuff, within their respective superclasses. The coarse prototype approach performs well for things classes but reduces stuff class performance due to the use of a single prototype for numerous samples. In contrast, the superclass prototype strategy achieves balanced results across both things and stuff, providing a +3.5 mIoU boost in things classes. This shows that creating prototypes by grouping classes with high visual similarity into superclasses is more effective for feature binding. Although the coarse prototype method enhances things classes, it causes a 1.5 mIoU decrease in stuff and a 0.2 mIoU drop in overall mIoU compared to the superclass prototype approach. This result shows that combining classes with low visual similarity into a single prototype provides suboptimal gains to things classes. Besides, improved prediction of things in driving scenes is beneficial, but poor prediction of stuff can disrupt long-term planning. Thus, the coarse prototype approach cannot be considered to produce ideal results.
Beam-wise Feature Distillation. We compare performance based on how feature distillation is applied and to which classes. Table 4 shows that limiting BFD to things classes provides better performance across both things and stuff compared to applying BFD to all classes. This likely occurs because stuff classes, with more points in the pitch direction, create class imbalance when BFD is applied. Additionally, point-wise feature matching in row 6 significantly reduces things classes performance compared to beam-wise distillation, as it fails to capture learning signals from dropped points, as noted in Sec. 3.4.
Overall Comparison. Applying superclass prototypes with beam-wise feature distillation only to things classes improves mIoU by 2.6 mIoU over the previous SoTA model, with a remarkable 4.8 mIoU increase in things classes performance.
4.4 Qualitative Results
Fig. 6 shows qualitative results of our method and baselines on the SemanticSTF validation set, using SemanticKITTI as the source domain. Correct predictions are marked in gray, errors are in red, and ground-truth classes are color-coded. Notably, our method outperforms SJLPD in segmenting things category, underscoring its intended gains for these critical classes. Specifically, in the dense fog scenario (a), our method accurately segments the motorcyclist class, while the original MinkowskiNet and SJ+LPD struggle to recognize it. This aligns with our quantitative result, with a 37.1 IoU gain for the motorcyclist class. Additionally, in rain and snow scenarios (c-d), our model excels in segmenting the person class, a critical class for autonomous driving due to its direct impact on safety.
5 Conclusion
In this study, we introduced our method, NTN, to address the limitations of existing DGLS-AW, particularly their poor performance on things classes. We identified two primary causes of this issue: (1) corruption of semantic-level features in things objects and (2) their vulnerability to local feature corruption caused by adverse weather conditions. To alleviate semantic-level feature corruption, we proposed a feature-binding approach that leverages superclass prototypes. To address the vulnerability of things classes to corruptions on local features, we introduced a beam-wise feature distillation loss that utilizes clean-weather beam-level features. Our methods establish a state-of-the-art on the SemanticKITTI-to-SemanticSTF and SemanticPOSS-to-SemanticSTF benchmarks, significantly boosting things classes performance.
Acknowledgements. This research was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the MSIP (RS-2025-00520207, RS-2023-00219019), KEIT grant funded by the Korean government (MOTIE) (No. 2022-0-00680, No. 2022-0-01045), the IITP grant funded by the Korean government (MSIT) (No. RS-2024-00457882, National AI Research Lab Project, No. 2021-0-02068 Artificial Intelligence Innovation Hub, RS-2019-II190075 Artificial Intelligence Graduate School Program (KAIST)), Samsung Electronics Co., Ltd (IO230508-06190-01) and SAMSUNG Research, Samsung Electronics Co., Ltd.
[…]
Supplementary Material
In this supplementary material, we provide additional details and results not included in the main paper due to space constraints. The content is organized in the following order:
  • Sec. S6. summarizes the rationale of our methods, Feature Binding and Beam-wise Feature Distillation.
  • Sec. S7. provides results across various weather conditions.
  • Sec. S8. offers qualitative results on the SemanticPOSS-to-SemanticSTF benchmark.
  • Sec. S9. provides experimental results of the SynLiDAR-to-SemanticSTF benchmark.
  • Sec. S10. provides discussions about the effectiveness and limitations of our methods.
  • Sec. S11 analyzes the mIoU drop for the bicyclist class.
  • Sec. S17. provides additional examples illustrating the comparison between clean and corrupted data.
S6 Rationale
In this section, we provide a detailed explanation of the rationale behind the two proposed methods, as introduced in the main paper.
Feature Binding. As outlined in the main paper, Feature Binding (FB) aims to prevent things objects from being mispredicted as stuff classes. Things objects are often misclassified as stuff due to semantic-level corruption caused by weather perturbations. In such conditions, accurately predicting fine-grained classes (e.g. person, motorcyclist) with LiDAR Semantic Segmentation models becomes highly challenging. FB mitigates this issue by constraining features to visually similar superclasses, reducing things-to-stuff mispredictions.
Beam-wise Feature Distillation. As discussed in the main paper, the goal of Beam-wise Feature Distillation (BFD) is to recover information lost due to missing points. BFD specifically addresses severe information loss in things objects by ensuring features from the augmented branch effectively capture things information from the clean branch. This approach aligns intact point patterns before the drop with collapsed point patterns after the drop, enabling efficient utilization of point pattern information.
Why Divide Corruptions into Semantic and Local Levels? This study categorizes corruptions based on the extent of point loss induced by weather corruption, which indicates potential degradation in data quality. A significant point loss can severely degrade semantic information. Thus, compensation strategies, such as Feature Binding, must be devised to mitigate this degradation at the semantic level. Even minimal point corruption can cause information loss. It disproportionately affects small-scale things objects. Beam-wise Feature Distillation is introduced to mitigate these corruptions.
Why FB Benefits things Classes More? Before applying our methodology, many misclassifications occurred between things and stuff classes. Thus, providing an additional discriminative signal via Feature Binding (FB) reduces misclassification between things and stuff. Second, we train the model to exploit the common point pattern in things classes through FB. This approach helps distinguish things from stuff because things share similar visual patterns, while stuff varies widely. Furthermore, as shown in Table 4, classwise prototypes do not improve performance. Therefore, our method’s effectiveness does not arise from alleviating class imbalance between things and stuff.
S7 Results on Specific Weather Conditions
SemanticSTF consists of four adverse weather conditions: dense fog, light fog, rain, and snow. We examine the performance under each weather condition. As shown in Table S5, on the SemanticKITTI-to-SemanticSTF benchmark, we observe that our model produces robust results regardless of the weather condition. Additionally, Table S6 shows the results on the SemanticPOSS-to-SemanticSTF benchmark. Our proposed model achieves the highest mIoU in most weather conditions, including significant improvements in rain and snow scenarios. Therefore, our model demonstrates consistent performance in most adverse weather scenarios across both benchmarks.
Method D-fog L-fog Rain Snow mIoU PolarMix [27] 29.7 25.0 28.6 25.6 27.2 PCL [29] 28.9 27.6 30.1 24.6 27.8 MMD [16] 30.4 28.1 32.8 25.2 29.1 PointDR [28] 31.3 29.7 31.9 26.2 29.8 DGLSS [12] 34.2 34.8 36.2 32.1 34.3 UniMix [31] 34.8 30.2 34.9 30.9 31.4 DGUIL [10] 36.3 34.5 35.5 33.3 34.8 SJ+LPD [20] 33.9 35.5 35.8 32.1 36.3 NTN (Ours) 35.3 35.1 35.7 32.4 38.9
Table S5: Performance comparison of different methods under varying weather conditions on SemanticKITTI-to-SemanticSTF benchmark. Bold indicates the best mIoU and underlined indicates the second-best performance.
Method D-fog L-fog Rain Snow mIoU PointDR [28] 26.2 30.1 50.1 43.2 34.7 DGLSS [12] 32.8 34.8 54.9 39.8 39.2 SJ+LPD [20] 25.4 30.6 38.2 40.6 38.3 NTN (Ours) 31.8 38.7 58.4 50.4 46.2
Table S6: Performance comparison of different methods under varying weather conditions on SemanticPOSS-to-SemanticSTF benchmark. Bold indicates the best mIoU and underlined indicates the second-best performance.
Figure S7: Qualitative results of our method on validation set of SemanticSTF. All models are trained on the train set of SemanticPOSS. Gray points indicate correct predictions, red points highlight errors, and ground-truth is shown with color-coded labels. Dashed circles highlight the predictions of things classes. Best viewed in color.
S8 Qualitative Results on the SemanticPOSS-to-SemanticSTF benchmark
Fig. S7 shows qualitative results of our method and baselines on the SemanticSTF validation set, using SemanticPOSS as the source domain. Correct predictions are marked in gray, errors are in red, and ground-truth classes are color-coded. Our method demonstrates significant improvements over the original MinkowskiNet and SJ+LPD in segmenting the things category, particularly for classes such as person and bicycle. In scenarios (a), (b), and (d), our method accurately segments the person class, while others fail to recognize it or generate incomplete segmentation results. These results align with our quantitative analysis, which reveals a 15.5 IoU improvement for the person class. Additionally, in the rain scenario (c), while the segmentation performance of bicycle is not yet optimal, our method still surpasses existing methods, providing better segmentation results. Accurate segmentation of these things classes is highly significant, as they directly impact the safety and reliability of autonomous driving.
S9 Experiment on SynLiDAR
Method car bi.cle mt.cle truck oth-v. pers. bi.clst mt.clst road parki. sidew. othe.g. build. fence veget. trunk terra. pole traf. mIoU Oracle 89.4 42.1 0.0 59.9 61.2 69.6 39.0 0.0 82.2 21.5 58.2 45.6 86.1 63.6 80.2 52.0 77.6 50.1 61.7 54.7 Source-only 27.1 3.0 0.6 15.8 0.1 25.2 1.8 5.6 23.9 0.3 14.6 0.6 36.3 19.9 37.9 17.9 41.8 9.5 2.3 15.0 Dropout [22] 28.0 3.0 1.4 9.6 0.0 17.1 0.8 0.7 34.2 6.8 30.5 1.1 35.5 19.1 42.3 17.6 36.0 14.0 2.8 15.2 Perturbation 27.1 2.3 2.3 16.0 0.1 23.7 1.2 4.0 27.0 3.6 16.2 0.8 29.2 16.7 35.3 18.3 17.9 5.1 2.4 15.2 PolarMix [27] 39.2 1.1 2.2 8.3 1.5 17.8 0.8 0.7 23.3 1.3 17.5 0.4 45.2 24.8 46.2 20.1 38.7 10.9 0.6 15.7 MMD [16] 25.5 2.3 2.1 13.2 0.7 22.1 1.4 7.5 30.8 0.4 17.6 0.4 30.9 19.7 37.6 19.3 43.5 9.9 2.6 15.1 PCL [29] 30.9 0.8 1.4 10.0 0.4 23.3 4.0 7.9 28.5 1.3 17.7 1.2 39.4 18.5 40.0 18.0 38.6 12.1 2.3 15.5 PointDR [28] 37.8 2.5 2.4 23.6 0.1 26.3 2.2 7.7 27.9 7.7 17.5 0.5 47.6 25.3 45.7 21.0 37.5 17.9 5.5 18.5 DGLSS [12] 47.9 2.9 3.4 17.4 1.1 28.0 2.4 7.3 28.8 10.2 18.1 0.2 48.9 25.3 46.5 21.4 45.2 17.9 4.9 19.8 UniMix [31] 65.4 0.1 3.9 16.9 5.3 32.3 2.0 19.3 52.1 5.0 27.3 3.0 49.4 20.3 58.5 22.7 23.2 26.1 20.9 23.4 DGUIL [10] 43.3 2.8 2.6 23.2 3.2 31.3 2.5 4.4 34.3 9.2 17.9 0.3 57.1 27.6 50.0 24.2 41.5 19.0 6.1 21.1 SJ+LPD [20] 39.0 2.5 2.5 22.3 0.3 27.0 1.8 4.0 36.1 10.3 19.0 1.0 50.6 24.5 45.1 23.2 34.1 21.9 7.2 19.6 NTN (Ours) 48.4 1.5 2.4 19.4 0.2 29.1 3.2 8.9 43.5 6.7 20.5 0.0 52.2 30.1 49.8 20.0 32.9 24.7 7.5 21.1 $\uparrow$ to SJ+LPD (+9.4) (-1.0) (-0.1) (-2.9) (-0.1) (+2.1) (+1.4) (+4.9) (+7.4) (-3.6) (+1.5) (-1.0) (+1.6) (+5.6) (+4.7) (-3.2) (-1.2) (+2.8) (+0.3) (+1.5)
Table S7: Comparison of methods on the SynLiDAR-to-SemanticSTF benchmark. Performance improvements of our method over SJ+LPD [20] are shown in the last row, with red text indicating things classes increments and green text for stuff.
We conduct additional experiments on the SynLiDAR-to-SemanticSTF benchmark. As shown in Table S7, our method improves the overall performance by 1.5 mIoU compared to the SJ+LPD model [20]. Although our proposed methods provide slight performance benefits, the improvements are not as significant as those observed in other benchmarks. SynLiDAR-to-SemanticSTF entangles both (1) the domain gap between synthetic and real data and (2) weather corruption. This is why our proposed methods provide marginal performance gain. Nevertheless, Table  S7 shows FB and BFD outperform the previous method, SJ+LPD [20], implying its effectiveness in representing real adverse weather. Note that UniMix [31] and DGUIL [10] have not been reproduced due to the lack of publicly available code.
S10 Discussions about Methods
In this section, we discuss both the effectiveness and limitations of our proposed modules on the SemanticKITTI-to-SemanticSTF benchmark.
Figure S8: Examples of the three levels of category classification: classes, superclasses, and coarse, in increasing order of granularity.
S10.1 Analysis on Feature Binding
Multi-Level Categories. We divide categories into three levels, as shown in Fig. S8. Classes are the smallest units, superclasses group semantically similar classes, and coarse categories consist of things and stuff. We perform a detailed analysis of these three levels using confusion matrices and qualitative results.
Figure S9: Confusion matrices for (a) superclasses and (b) things and stuff after adding Feature Binding. The numbers in brackets indicate the increase or decrease compared to SJ+LPD [20]. FB effectively reduces misclassification between things and stuff.
Effect of FB for Relieving Confusion. As mentioned in Sec. S6, Feature Binding is a module designed to prevent things from being misclassified as stuff when their semantic information is partially degraded due to corruption caused by adverse weather. To examine the effect of FB, we conduct an analysis at the superclass level and the coarse level, as shown in Fig. S9. When we compare the performance with the previous method [20], we find that misclassification ratio of things as stuff decreases by 2.34%. At the superclass level, we observe that the tendency to misclassify things as natural categories such as vegetation is greatly reduced. In particular, for the person category, true positive prediction ratio increases by 9.22% compared to before, and misclassification ratio as the natural category decreases by 5.74%. Thus, we confirm that FB helps to learn the semantic differences between things and stuff by continuously providing hierarchical semantic information.
Figure S10: Class-wise confusion matrices of (a) SJ+LPD [20] and (b) Ours on SemanticKITTI-to-SemanticSTF benchmark.
Figure S11: Qualitative results of NTN (FB + BFD). Gray is true prediction, and Red is false prediction. Other colors mean true predictions for specific classes.
Feature Distillation Types Performance $\mathcal{B}$-all. $\mathcal{B}$-wise. (Ours) Things Stuff mIoU 35.3 43.7 38.9 36.2 42.5 38.9
Table S8: Comparison for feature distillation settings between average on all beams and seperate beam. Experiments are done on SemanticKITTI-to-SemanticSTF benchmark.
S10.2 Analysis on Beam-wise Feature Distillation.
Challenges of BFD with Disrupted Point Patterns. As in Fig. S10, misprediction ratio from things-to-stuff slightly increases in bicycle and bicyclist classes. This is because BFD tends to effectively utilize the point patterns before point drop, as mentioned in Sec. S6.
The learning mechanism of BFD enables the LSS model to perform better on objects where the clean branch’s point pattern is well preserved during test time. This explains why our method struggles with classes like bicycle and bicyclist, which inherently have few points and are prone to severe point pattern corruptions. In such cases, BFD struggles to utilize informations of local point pattern from clean data branch. Conversely, significant performance improvements are observed for classes like motorcycle and motorcyclist, which have clearer shapes. As shown in Fig. S11, the method with BFD demonstrates superior predictions for bicycles with reasonably preserved point patterns, as observed in row 1. Also, the motorcycle in row 2, which maintains its point pattern, achieves improved performance. In summary, confusion on classes like bicycle arises as a side effect of improving performance by preserving things object information without bias toward weather-corrupted data.
Ablation Study on Types of Feature Distillation. As shown in Table S8, distillation after averaging across all beams ($\mathcal{B}$-all.) achieves the same mIoU as our method, which performs beam-wise averaging before distillation ($\mathcal{B}$-wise.). However, our method outperforms $\mathcal{B}$-all. in things objects. This demonstrates that defining beam-wise local regions and performing distillation accordingly better compensates for information loss caused by point missing in things objects.
Figure S12: Qualitative results on bicyclist class for PolarMix [27], PointDR [28], SJ+LPD [20] and Ours.
Dataset Sem.KITTI; w. Other Sim. Dataset nuScenes; real rain nuScenes-C; snow, fog, wet ground
Method Things Stuff All Method Things Stuff All Heavy Moderate Light Things Stuff All
Source Only 17.7 38.9 26.6 Source Only 34.2 67.0 48.6 47.5 52.5 55.1 39.0 68.0 51.7
[8] + [9] 22.9 42.8 31.3 SJ+LPD 33.2 66.6 47.8 49.0 53.6 55.7 40.7 68.2 52.7
+ NTN (Ours) 23.6 43.9 32.2 + NTN (Ours) 35.4 67.8 49.6 48.6 54.0 56.6 53.1 68.5 53.1
$\uparrow$ to [8] + [9] (+0.7) (+1.1) (+0.9) $\uparrow$ to SJ+LPD (+2.2) (+1.2) (+1.8) (+2.2) (+1.9) (+1.7) (+12.4) (+0.3) (+0.4)
Table S9: Comparison on various benchmarks and scenarios.
Method Superclass mIoU Total Coarse Things: person, vehicle, traffic element 33.7 ($\uparrow$2.4) 38.1 ($\uparrow$1.8) Stuff: pavement, natural, structure 44.0 ($\uparrow$0.9) GPT-o1 Person: person, bi.clst, mt.clst 39.1 ($\uparrow$13.4) 38.3 ($\uparrow$2.0) Vehicle: car, bi.cle, mt.cle, truck, oth-v. 33.5 ($\uparrow$3.4) Ground: road, parki., sidew., othe.g. 30.0 ($\downarrow$0.8) Nature: trunk, veget., terra. 48.6 ($\uparrow$0.4) Construction: build., fence 57.3 ($\downarrow$6.8) Traffic Object: pole, traf. 31.5 ($\downarrow$2.0) DeepSeek-V3 Person: person, bi.clst, mt.clst 30.0 ($\uparrow$4.4) 39.4 ($\uparrow$3.1) Vehicle: car, bi.cle, mt.cle, truck, oth-v. 37.3 ($\uparrow$7.2) Road-infra.: road, parki., sidew., othe.g. 32.1 ($\uparrow$1.3) Nature: veget., terra. 56.6 ($\uparrow$2.9) Man-made: build., fence, pole, traf. 49.1 ($\uparrow$0.4) Miscell.: trunk 35.1 ($\downarrow$2.1)
Table S10: Comparison of different superclasses on SemanticKITTI-to-SemanticSTF (improvements over SJ+LPD in brackets).
S11 Analysis on mIoU Drop Bicyclist & Fence Classes.
Bicyclist. We found out that both the predictions from SJ+LPD [20] and our method accurately detected the bicyclist within a single scan. As shown in the third and fourth visualizations of Fig. S12, the performance difference between SJ+LPD and our method is minimal, amounting to only a few points. Furthermore, the mispredicted points were all classified as a person, indicating that the slight performance drop does not significantly impact safety-critical predictions.
We verified the results for this object using PointDR and PolarMix, which are reproducible due to publicly available codes. As illustrated in the first and second visualizations of Fig. S12, both PolarMix [27] and PointDR [28] failed to provide accurate predictions for most bicycle points. In PolarMix, some points are predicted as bicyclist, but many are misclassified as motorcyclist. As motorcyclist generally moves faster than bicyclist or person, such mispredictions could lead to significant risks in ensuring safe driving. For PointDR, none of the points are predicted as person, bicyclist, or motorcyclist. This demonstrates that previous methods fail to predict even within the person superclass, highlighting the effectiveness of our method in safety-critical driving scenarios.
Fence. Unlike other thin traffic elements, fence varies from thin to large structures. Due to this structural diversity, large fences receive confusing signals from the distinct-shaped superclass by FB. As a result, fence had lower IoU and increased confusion with building and vegetation as in Fig. S10.
[…]
The performance gap arises from differences in car class annotation between SemanticKITTI and SemanticPOSS. SemanticKITTI-trained model benefits from FB by learning diverse vehicle classes separately, thus effective in SemanticSTF. In contrast, SemanticPOSS groups car, bus, and truck into one class, limiting feature diversity and lowering performance. For optimal performance of FB, pre-defined classes should be as fine-grained as possible.
S13 Related Works on Subclass- or Prototype-based Methods
Subclass- or Prototype-based regularization has frequently been employed to tackle label-efficiency problems, where limited or imbalanced annotations lead to suboptimal segmentation. Pixel-to-Prototype Contrast [6] addresses weakly supervised semantic segmentation by aligning pixel embeddings with class prototypes, refining noisy pseudo masks generated from image-level labels. Similarly, Prototypical Contrastive Network [18] focuses on highly imbalanced aerial segmentation by learning a single foreground prototype and pushing away hard-negative background features, thus emphasizing minority classes. Unbiased Subclass Regularization [7] aims to mitigate semi-supervised class imbalance by splitting overrepresented classes into smaller clusters, forming class-balanced subclasses. Despite sharing the general goal of improving segmentation via prototype or class-based groupings, our work diverges in both methodology and application. Instead of subdividing a single class or maintaining per-class prototypes, we merge multiple classes into higher-level superclasses (e.g., things vs. stuff), which better addresses the broad semantic gap under adverse weather in LiDAR data. Moreover, prior methods largely concentrate on 2D image tasks or semi-/weakly supervised settings, whereas we focus on single-domain generalization for 3D LiDAR segmentation, emphasizing robustness against severe weather-induced corruptions.
S14 Additional Experiments on Different Weather Simulations.
We used simulation methods [8, 9] to generate occlusion-induced points, improving performance (Table S9) and confirming robustness to other weather simulations. Similar to Table 4, the performance gain was higher in things (+0.7 mIoU), highlighting our method’s safety benefits.
S15 Additional Experiments on Various Datasets
We tested our method on two settings: (1) nuScenes clean train set-to-rainy splits, and (2) nuScenes train set-to-nuScenes-C. As shown in Table S9, our method consistently surpasses SJ+LPD [20] across diverse datasets, particularly for things.
S16 Additional Experiments for Different Superclass.
[…]
S17 Additional Visual Comparisons Between Clean and Corrupted Data
We provide additional comparisons between clean and corrupted data, as shown in Fig. S13. It illustrates point patterns of things objects in clean and weather-induced corrupted data. As mentioned in Sec. 1 and Sec. 3 in the main paper, things objects in clean weather have well-defined shapes and smooth boundaries, whereas in adverse weather, they have blurred shapes and irregular boundaries. This proves that things classes are more vulnerable to such noise or point loss, making accurate predictions more difficult.
Figure S13: Examples of clean and corrupted data for various things classes. In clean weather, objects have well-constructed shapes with dense point clouds. In contrast, objects in adverse weather have blurred shapes with significant point loss.
Figure S14: Qualitative results of failure cases on the validation set of SemanticSTF. Ground-truth and our prediction results are shown with color-coded labels. For comparison, gray points indicate correct predictions and red points highlight errors.
S18 Failure Cases
Fig. S14 illustrates cases where our method fails in prediction. Errors remain, particularly at very close distances to vehicles, as shown in (a) and (e). As demonstrated in (a), performance improvements were limited for the car class due to extreme sparsity from occlusion by droplets. For the person class, (c) shows yet many incorrect predictions remain in extremely noisy conditions. Thin objects like pole and traffic-sign are often misclassified as vegetation, as seen in cases (b), (d), (e), (f) and (h). Errors also occur in classes with varying object sizes, such as fences in examples (g) and (h). As discussed in Sec. S11, this issue arises from defining FB superclasses manually.

### v2

@@
 This difference arises because things classes are generally small and dispersed, leaving them vulnerable to missing points that can distort their structure.
-For example, person class, belonging to the things category, have dense and clear shapes under clean weather, as shown on the far left of the figure.
+For example, person class, belonging to the things category, have dense and clear shapes under clean weather, as shown on the far left of the Fig.
+2 (a) and (b).
 Consequently, under adverse weather conditions, the person and car points are fewer in number and have obscure shapes.
@@
 PointDR [28] generates a prototype in the clean weather branch and multiplies corrupted branch features by this prototype matrix, utilizing clean data insights.
-DGUIL [10] addresses point uncertainty under domain shifts by by improving point features across varying distributions in uncertain point clouds.
+DGUIL [10] addresses point uncertainty under domain shifts by improving point features across varying distributions in uncertain point clouds.
 However, existing weather simulation and domain generalization methods focus on replicating weather conditions and do not address precise recognition of key things classes under weather-induced corruption in driving environments.
@@
 3.1, we first introduce a concept of Feature Binding loss (FB).
-The purpose of FB loss is to “bind” the output features of LSS model, preventing misprediction of things objects as stuff when local corruption severely affects semantic-level features.
+The purpose of FB loss is to “bind” the output features of LSS model, preventing misprediction of things objects as stuff when weather corruption severely affects semantic-level features.
 To achieve this, we create prototypes for superclasses by grouping visually similar pre-defined classes during LSS training.
 Tab.
-S10 summarizes the superclass categories and their corresponding classes.
+1 summarizes the superclass categories and their corresponding classes.
 Six superclasses are defined for SemanticKITTI [3], while four superclasses are defined for SemanticPOSS [19].
@@
 The SemanticPOSS dataset contains a higher proportion of things classes at 28.9%, compared to SemanticSTF, and is known for its diversity.
-Consequently, the model has already learned relatively rich features, resulting in less performance improvement for things classes compared to SemanticKITTI.
+Consequently, the model has already learned relatively rich features, resulting in less performance improvement for things classes compared to SemanticKITTI-to-SemanticSTF benchmark.
 Figure 6: Qualitative results of our method on validation set of SemanticSTF.
@@
 Although the coarse prototype method enhances things classes, it causes a 1.5 mIoU decrease in stuff and a 0.2 mIoU drop in overall mIoU compared to the superclass prototype approach.
-This result shows that combining classes with low visual similarity into a single prototype provides suboptimal gains to things classes.
+This result shows that combining classes with low visual similarity into a single prototype provides suboptimal gains.
 Besides, improved prediction of things in driving scenes is beneficial, but poor prediction of stuff can disrupt long-term planning.
@@
   • Sec. S10. provides discussions about the effectiveness and limitations of our methods.
-  • Sec. S11 analyzes the mIoU drop for the bicyclist class.
+  • Sec. S11 analyzes the mIoU drop for the bicyclist and fence class.
+  • Sec. S12. provides discussions about the performance degradation of car class on SemanticPOSS-to-SemanticSTF benchmark.
+  • Sec. S13. provides related works on subclass- or prototype- based methods.
+  • Sec. S14. provides additional experiments on different weather simulations.
+  • Sec. S15. provides additional experiments on various datasets.
+  • Sec. S16. provides additional experiments for different superclass criterions.
   • Sec. S17. provides additional examples illustrating the comparison between clean and corrupted data.
+  • Sec. S18. provides failure cases of our methods.
 S6 Rationale